# Parallel Corpora for Machine Translation in Low-Resource Indic Languages: A Comprehensive Review

### v1

@@ 4.1 Expanding Coverage for Low-Resource and Dialectal Variants
Many Indic languages, such as Santali, Bodo, Manipuri, and Konkani, remain underrepresented in existing parallel corpora. Most datasets focus on high-resource languages like Hindi, Bengali, and Tamil, creating an imbalance that hinders the development of MT models for low-resource languages  [86]. Future efforts should prioritize the collection of bilingual and multilingual parallel data from vernacular media, government archives, social media, and oral histories  [78]. Crowdsourcing initiatives and community-driven data curation can further help improve linguistic diversity and increase the representation of marginalized languages in NLP applications.
@@ 4.2 Enhancing Domain Adaptation and Contextual Alignment
Most existing Indic parallel corpora are domain-specific, with a strong bias toward news, religious texts, and government documents. This limits their applicability in scientific, medical, legal, and conversational domains  [9]. To improve cross-domain generalization, future work should focus on constructing multi-domain parallel corpora and training domain-adaptive MT models  [87]. Furthermore, context-aware alignment techniques, such as document-level parallel corpora and sentence embedding-based alignment, can enhance semantic consistency and translation fluency across diverse textual genres.
@@ 4.3 Leveraging Multimodal and Code-Switched Parallel Data
Multimodal MT, which involves image-text and speech-text parallel corpora, is becoming increasingly relevant for Indic languages. Initial datasets like Hindi Visual Genome and Bengali Visual Genome demonstrate the potential of multimodal learning, but larger, more diverse datasets are needed to improve multimodal translation systems  [57]. Similarly, code-switching is prevalent in Hindi-English, Bengali-English, and Tamil-English interactions, yet parallel code-switched corpora remain scarce. Expanding multimodal and code-switched datasets will enhance MT models’ performance in real-world multilingual communication and improve their ability to handle informal language  [88].
@@ 4.4 Automatic Data Generation
Given the scarcity of human-annotated parallel corpora, automatic data generation techniques such as back-translation, synthetic data augmentation, and parallel data mining have gained prominence  [89]. Back-translation, where monolingual target-language data is translated into the source language using pre-trained models, has been widely used to augment data for low-resource Indic languages  [90]. Similarly, parallel sentence mining techniques, such as LASER have been applied to extract sentence-aligned parallel data from large-scale web corpora. Additionally, zero-shot learning and self-supervised approaches can further help bootstrap translation models for languages with minimal parallel data. Future work should focus on refining these techniques to improve alignment accuracy and reduce noise in automatically generated corpora.
@@ Appendix A Appendix
Itihasa Text-to-Text 93K English-Sanskrit Itihasa Parallel Corpus
BUET English-Bangla Text-to-Text 2.7M English-Bangla BUET English-Bangla Corpus
@@
IndoWordNet Text-to-Text 6.3M 18 Indic languages IndoWordNet Parallel Corpus
@@
Uka Tarsadia Text-to-Text 65K English-Gujarati Uka Tarsadia Corpus
@@
BhasaAnuvaad Speech¡-¿Text 44,400+ hours 13 indic languages BhasaAnuvaad
Mann ki Baat Speech¡-¿Text N/A Hindi Mann ki Baat
IndicTTS Speech¡-¿Text 100+ hours/language 7 Indic languages IndicTTS
@@
Hindi Visual Genome Multimodal  31K Hindi-English Hindi Visual Genome
Bengali Visual Genome Multimodal 29K Bengali-English Bengali Visual Genome
Malayalam Visual Genome Multimodal 29K Malayalam-English Malayalam Visual Genome
Table 1: Overview of major parallel corpora available for Indic languages, including key characteristics such as language coverage, dataset size, and translation type.
@@
Table 2: Overview of the diverse scripts used in Indic languages, categorized by geographical regions and linguistic families across South Asia. It highlights the rich linguistic heritage and the varying script traditions prevalent in the region
Fig. 3: Indic Parallel Corpora: Data Size Distribution Across Languages – A stacked bar chart showing the availability of sentence pairs for different language pairs across multiple corpora. Only few languages like Hindi, Tamil and Bengali are well represented.

### v2

@@ 4.1 Expanding Coverage for Low-Resource and Dialectal Variants
Many Indic languages, such as Santali, Bodo, Manipuri, and Konkani, remain underrepresented in existing parallel corpora. Most datasets focus on high-resource languages like Hindi, Bengali, and Tamil, creating an imbalance that hinders the development of MT models for low-resource languages  Lupascu et al. (2025). Future efforts should prioritize the collection of bilingual and multilingual parallel data from vernacular media, government archives, social media, and oral histories  Guzmán et al. (2019). Crowdsourcing initiatives and community-driven data curation can further help improve linguistic diversity and increase the representation of marginalized languages in NLP applications.
@@ 4.2 Enhancing Domain Adaptation and Contextual Alignment
Most existing Indic parallel corpora are domain-specific, with a strong bias toward news, religious texts, and government documents. This limits their applicability in scientific, medical, legal, and conversational domains  Hu et al. (2024b). To improve cross-domain generalization, future work should focus on constructing multi-domain parallel corpora and training domain-adaptive MT models  Dong et al. (2025). Furthermore, context-aware alignment techniques, such as document-level parallel corpora and sentence embedding-based alignment, can enhance semantic consistency and translation fluency across diverse textual genres.
@@ 4.3 Leveraging Multimodal and Code-Switched Parallel Data
Multimodal MT, which involves image-text and speech-text parallel corpora, is becoming increasingly relevant for Indic languages. Initial datasets like Hindi Visual Genome and Bengali Visual Genome demonstrate the potential of multimodal learning, but larger, more diverse datasets are needed to improve multimodal translation systems  Sen et al. (2021). Similarly, code-switching is prevalent in Hindi-English, Bengali-English, and Tamil-English interactions, yet parallel code-switched corpora remain scarce. Expanding multimodal and code-switched datasets will enhance MT models’ performance in real-world multilingual communication and improve their ability to handle informal language  Winata et al. (2021b).
@@ 4.4 Automatic Data Generation
Given the scarcity of human-annotated parallel corpora, automatic data generation techniques such as back-translation, synthetic data augmentation, and parallel data mining have gained prominence  Shu et al. (2024). Back-translation, where monolingual target-language data is translated into the source language using pre-trained models, has been widely used to augment data for low-resource Indic languages  Sennrich et al. (2016). Similarly, parallel sentence mining techniques, such as LASER have been applied to extract sentence-aligned parallel data from large-scale web corpora. Additionally, zero-shot learning and self-supervised approaches can further help bootstrap translation models for languages with minimal parallel data. Future work should focus on refining these techniques to improve alignment accuracy and reduce noise in automatically generated corpora.
@@ Appendix A Appendix
Itihasa Text-to-Text 93K English-Sanskrit Itihasa Corpus
BUET Eng-Bn Corpus Text-to-Text 2.7M English-Bangla BUET Corpus
@@
IndoWordNet Text-to-Text 6.3M 18 Indic languages IndoWordNet Corpus
@@
Uka Tarsadia Text-to-Text 65K English-Gujarati Uka Tarsadia
@@
BhasaAnuvaad Speech<->Text 44,400+ hrs 13 Indic languages BhasaAnuvaad
Mann ki Baat Speech<->Text N/A Hindi Mann ki Baat
IndicTTS Speech<->Text 100+ hrs/language 7 Indic languages IndicTTS
@@
Hi Visual Genome Multimodal  31K Hindi-English Hindi Visual Genome
Bn Visual Genome Multimodal 29K Bengali-English Bengali Visual Genome
Ml Visual Genome Multimodal 29K Malayalam-English Malayalam Visual Genome
Table 1: Comprehensive overview of major parallel corpora available for Indic languages, spanning various modalities including text-to-text, speech-to-text, code-switched, and multimodal datasets. The table highlights the corpus type, number of sentence pairs or duration (where applicable), supported language pairs (with a focus on English-Indic and intra-Indic combinations), and links to official sources for access. These resources play a vital role in enabling research in machine translation, multilingual NLP, and low-resource language processing across the Indic language spectrum.
@@
Table 2: Comprehensive overview of the diverse scripts used across Indic languages, categorized by associated languages, geographic regions, and their respective script families. This table, referenced in Section 1.2.2, highlights both modern and historical scripts, including Brahmic-derived scripts (e.g., Devanagari, Tamil, Bengali), Perso-Arabic adaptations (e.g., Urdu, Kashmiri), and lesser-known indigenous scripts (e.g., Meitei Mayek, Lepcha, Chakma). The representation illustrates the linguistic diversity and orthographic complexity of the Indian subcontinent—factors that critically affect text normalization, OCR development, and
Figure 3: Indic Parallel Corpora: Data Size Distribution Across Languages. A stacked bar chart showing the number of sentence pairs (in millions) for various Indic-English language pairs across major corpora. Languages like Hindi, Tamil, and Bengali are well-resourced, while others such as Assamese and Odia have significantly less data. This highlights the data imbalance in Indic NLP and the need for better resource coverage.